# Survey of Video Diffusion Models: Foundations, Implementations, and Applications

### v1

@@ Abstract
Recent advances in diffusion models have revolutionized video generation, offering superior temporal consistency and visual quality compared to traditional generative adversarial networks-based approaches. While this emerging field shows tremendous promise in applications, it faces significant challenges in motion consistency, computational efficiency, and ethical considerations. This survey provides a comprehensive review of diffusion-based video generation, examining its evolution, technical foundations, and practical applications. We present a systematic taxonomy of current methodologies, analyze architectural innovations and optimization strategies, and investigate applications across low-level vision tasks such as denoising and super-resolution. Additionally, we explore the synergies between diffusion-based video generation and related domains, including video representation learning, question answering, and retrieval. Compared to the existing surveys (Lei et al., 2024a; b; Melnik et al., 2024; Cao et al., 2023; Xing et al., 2024c) which focus on specific aspects of video generation, such as human video synthesis (Lei et al., 2024a) or long-form content generation (Lei et al., 2024b), our work provides a broader, more updated, and more fine-grained perspective on diffusion-based approaches with a special section for evaluation metrics, industry solutions, and training engineering techniques in video generation. This survey serves as a foundational resource for researchers and practitioners working at the intersection of diffusion models and video generation, providing insights into both the theoretical frameworks and practical implementations that drive this rapidly evolving field. A structured list of related works involved in this survey is also available on GitHub.
@@ 1 Introduction
The rest of the survey is organized as follows: Section 2 lays the groundwork by exploring foundational concepts, including video generation paradigms such as GAN-based models, auto-regressive models, and diffusion models. Section 3 focuses on implementation, addressing practical aspects such as datasets, training engineering techniques, and evaluation metrics, along with benchmarking findings to highlight model performance. Section 4 covers diverse applications, including conditional generation tasks, enhancement methods such as video denoising, inpainting, interpolation, extrapolation, and super-resolution, as well as personalization, consistency, long-video generation, and emerging 3D-aware diffusion models. Finally, Section 5 discusses the benefits of diffusion-based video generation to other domains, such as video representation learning, retrieval, QA, and 3D and 4D generation, emphasizing its broader impact in advancing related fields. This structure provides a comprehensive and cohesive understanding of diffusion-based video generation, from foundational principles to advanced applications.
2 Foundation
2.1 Video generative paradigm
@@ 2.1.1 GAN video models
GANs consist of a generator and a discriminator. The generator aims to produce realistic samples, while the discriminator seeks to differentiate between the generated samples and the ground truth samples. For video generation, GANs have been extended to model temporal consistency and generate realistic video frames. For example, TGAN (Saito et al., 2017) uses two generators, with a temporal generator creating motion features and an image generator producing video frames. TGAN-v2 (Saito et al., 2020) improves efficiency by generating videos at multiple temporal resolutions using a cascade of generator modules. MoCoGAN (Tulyakov et al., 2018) separates latent space into motion and content parts, predicting motion vectors auto-regressively, while DVD-GAN (Clark et al., 2019) generates all frames in parallel and conditions on class labels. MoCoGAN-HD (Tian et al., 2020) disentangles content and motion differently, with a motion generator predicting a latent motion trajectory. DIGAN (Yu et al., 2021) further leverages implicit neural representations to efficiently generate longer video sequences with a motion discriminator focused on frame pairs rather than entire sequences. StyleGAN2 (Karras et al., 2020) introduced a highly expressive and controllable latent space for photorealistic image synthesis, laying a foundational framework for video synthesis through sparse motion cues and efficient control mechanisms. Building on this, StyleGAN-(Skorokhodov et al., 2022) advanced the field by modelling videos as continuous-time signals, improving temporal coherence while supporting high-resolution and long-duration generation. Extending these developments, StyleInv (Wang et al., 2023i)) proposed a non-autoregressive GAN inversion network, achieving smoother motion transitions and enabling style transfer through fine-tuning. AniFaceGAN (Wu et al., 2022c) adapted StyleGAN2 into an animatable 3D-aware GAN, enabling multiview-consistent face animation generation, particularly for face-centric applications. Recently, SIDGAN (Muaz et al., 2023) tackled high-resolution dubbed video generation by introducing shift-invariant learning and a pyramid-based decoder, enabling accurate lip-syncing and high-quality textures even under challenging conditions. Further advancing lip-sync technology, StyleLipSync (Ki & Min, 2023) leveraged pre-trained StyleGAN to generate temporally consistent lip-synced videos with pose-aware masking and moving-average-based latent smoothing, enabling high-fidelity results for arbitrary identities. Details about the GAN application on video generation can be found  (Aldausari et al., 2022).
@@ 2.1.3 Video diffusion models
GAN-based and Auto-Regressive models often face challenges in maintaining temporal consistency and computational efficiency when generating videos. While GANs struggle with maintaining temporal consistency and often produce unstable or inconsistent frames, diffusion models excel at generating smooth, coherent sequences, particularly for long videos, because their iterative denoising process ensures frame-to-frame consistency. Also, despite Auto-Regressive model can handle both continuous and discrete data, it normally requires high computational cost. In terms of frame quality, diffusion models generate more detailed and higher-resolution outputs compared to GANs, which are prone to artifacts from adversarial training, and autoregressive models, which face resolution limitations due to the compounding of errors in sequential prediction[cite]. Diffusion models also offer superior controllability and local editing capabilities because they can refine specific parts of the video during the denoising process without affecting other regions while GANs and autoregressive methods generate frames in a more holistic or sequential manner[cite].
@@ 2.1.4 Auto-regressive video diffusion models
Some recent works further exploit auto-regressive framework in the context of video generation with diffusion models (Weng et al., 2024; Jin et al., 2024; Xie et al., 2024a). ART•V (Weng et al., 2024) first utilized autoregressive diffusion to generate frames sequentially, employing masked diffusion to enhance coherence and reduce inconsistencies. Subsequently, Pyramidal Flow Matching  (Jin et al., 2024) introduced spatial and temporal pyramids to optimize autoregressive generation, reducing computational costs and improving scalability. Progressive Autoregressive Video Diffusion Models  (Xie et al., 2024a) refined this by using a progressive noise schedule, enabling smoother transitions in long videos. CausVid  (Yin et al., 2024) advanced the field by converting bidirectional diffusion models into causal autoregressive generators through asymmetric distillation, achieving real-time frame-by-frame generation with low latency and reduced error accumulation.
2.2 Learning foundation
@@ 2.2.7 Training-free methods
Text2Video-Zero (Khachatryan et al., 2023) introduced cross-frame attention based on pretrained text-to-image diffusion model. FateZero (Qi et al., 2023) enhanced this through intermediate attention maps, and Free-Bloom (Huang et al., 2023a) combined GPT-3 with image diffusion models for improved semantic coherence. PEEKABOO (Jain et al., 2024b) enabled spatio-temporal control through masked diffusion using a latent diffusion model-based video generation model, while ControlVideo (Zhang et al., 2023e) extended ControlNet with an interleaved-frame smoother to generate consistent controllable and structurally smooth video. DreamVideo-2 (Wei et al., 2024b) proposed zero-shot subject-driven video customization with precise motion control, enabling subject-preserving and motion-specific video editing. Following this, VideoElevator (Zhang et al., 2024f) encapsulates T2V to enhance temporal consistency and harnesses T2I to provide more faithful details. Recent innovations include GPT4Motion (Lv et al., 2024)’s combination of GPT-4, Blender, and text-to-image diffusion models for physics-aware generation, AnyV2V (Ku et al., 2024)’s universal editing framework built on an image editing model combined with an existing image-to-video generation model to generate the edited video through temporal feature injection. Video Custom Diffusion (Magic-Me) (Ma et al., 2024e) built a training-free 3D Gaussian Noise Prior for video frame initialization, reconstructing inter-frame correlation to achieve identity-preserved generation.
@@ 2.2.8 Token learning
2.3 Guidance
@@ 2.3.1 Classifier guidance
Algorithm for classifier-guided DDPM sampling. In the DDPM framework, classifier guidance can be incorporated by modifying the reverse diffusion process. The update at each timestep involves sampling from a Gaussian distribution with a mean adjusted by the classifier gradients. Detailed workflow can be referred to Algorithm algorithm 1.
Algorithm 1 Classifier-guided DDPM sampling, given a diffusion model $(\mu_{\theta}(x_{t}),\Sigma_{\theta}(x_{t}))$, classifier $p_{\phi}(y|x_{t})$, and gradient scale $s$.(Dhariwal & Nichol, 2021)
@@
Algorithm for classifier-guided DDIM sampling. For the DDIM framework, the deterministic nature of the sampling process requires a slightly different approach. The classifier gradient is used to modify the epsilon prediction, which guides the reverse sampling process in DDIM. Deatailed workflow can be referred to Algorithm algorithm 2.
@@ 2.3.2 Classifier-free guidance
Algorithm for Classifier-Free Guidance describes the sampling procedure for classifier-free guidance, which adjusts the score estimates directly within the diffusion model without requiring any classifier gradients. Algorithms algorithm 3 and algorithm 4 describe training and sampling with classifier-free guidance in detail.
@@
3:     $(x,y)\sim p(x,y)$triangleright Sample data with conditioning information from the dataset
4:     $y\leftarrow\emptyset$ with probability $p_{\mathrm{uncond}}$triangleright Randomly discard conditioning to train unconditionally
5:     $\lambda\sim p(\lambda)$triangleright Sample log SNR value
@@
7:     $x_{\lambda}=\alpha_{\lambda}x+\sigma_{\lambda}\epsilon$triangleright Corrupt data to the sampled log SNR value
8:     Take gradient step on $\nabla_{\theta}\|\epsilon_{\theta}(x_{\lambda},y)-\epsilon\|^{2}$triangleright Optimize the denoising model
@@
3:$\lambda_{1},\ldots,\lambda_{T}$: increasing log SNR sequence with $\lambda_{1}=\lambda_{\text{min}},\lambda_{T}=\lambda_{\text{max}}$
4:$x_{0}\sim\mathcal{N}(0,I)$
@@
6:     $\tilde{\epsilon}_{t-1}\leftarrow(1+w)\epsilon_{\theta}(x_{t-1})-w\epsilon_{\theta}(x_{t-1})$triangleright Form the classifier-free guided score at log SNR $\lambda_{t-1}$
7:     $x_{t}\leftarrow\frac{(x_{t-1}-\alpha_{\lambda_{t-1}}\tilde{\epsilon}_{t-1})}{\alpha_{\lambda_{t-1}}}$triangleright Sampling step (could be replaced by another sampler, e.g., DDIM)
@@ 2.4.2 Optical-flow-based diffusion models
Another distinct approach within diffusion models leverages optical flow to maintain temporal coherence between frames. Optical flow is a common approach to represent motion by estimating the displacement field between consecutive frames. Early approaches treated it as an optimization problem using handcrafted features to maximize visual similarity(Horn & Schunck, 1981; Black & Anandan, 1993; Bruhn et al., 2005; Sun et al., 2014). Incorporating deep learning-based methods revolutionized this field. FlowNet(Dosovitskiy et al., 2015) pioneered end-to-end optical flow estimation, demonstrating the potential of deep learning in this domain. Subsequent advancements on improved architectures and synthetic datasets further promoted the progress of optical flow estimation(Ilg et al., 2017; Ranjan & Black, 2017; Sun et al., 2018; 2021; Hui et al., 2018; 2020; Yang & Ramanan, 2019). RAFT(Teed & Deng, 2020) included iterative refinement with correlation volumes, significantly boosting performance, while FlowFormer(Huang et al., 2022b; Shi et al., 2023c) applied attention mechanisms within optical flow usage. VideoFlow(Shi et al., 2023b) extended optical flow to video generation scenarios and utilized temporal information across multiple frames, achieving high accuracy. (Hu et al., 2023)propose a Dynamic Multi-scale Voxel Flow Network (DMVFN) to explicitly model the complex motion cues of diverse scales between adjacent video frames by dynamic optical flow estimation to generate videos at lower computational costs. Also, Motion-I2V(Shi et al., 2024) further refined the motion modelling by a diffusion-based motion field predictor and motion-augmented temporal attention to generate more consistent videos even in the presence of large motion and viewpoint variation. By conditioning the optical flow between a previous and a subsequent frame, this framework achieves realistic motion continuity, which is critical for applications in video generation and motion synthesis. Additionally, recent advancements explore upgrading pretrained image diffusion models to video generation by temporally warping input noise. How I Warped Your Noise (Chang et al., 2024) proposed a novel noise representation called integral noise ($\int$-noise), preserving temporal correlations through an optical-flow-based noise warping algorithm. This foundational approach mitigated issues like high-frequency flickering and texture-sticking artifacts, laying the groundwork for temporally coherent video diffusion. Building upon this, Infinite-Resolution Integral Noise Warping (Deng et al., 2025) method introduces a training-free approach where temporal correlations in noise are preserved by warping input noise using integral representations, enabling consistent frame generation while reducing computational overhead. These diffusion model frameworks demonstrate the versatility and adaptability of diffusion processes across different domains, from high-resolution static images to complex, temporally coherent videos. As shown in Figure 3, Go-with-the-Flow (Burgert et al., 2025) combined optical flow extraction from training videos with real-time noise warping and fine-tuning video diffusion models on paired warped noise and video data to enable robust and motion-controllable generation. Together, these methods demonstrate the progression from temporal correlation preservation to advanced motion control, showcasing the versatility and adaptability of diffusion models in video generation.

### v2

@@ Abstract
Recent advances in diffusion models have revolutionized video generation, offering superior temporal consistency and visual quality compared to traditional generative adversarial networks-based approaches. While this emerging field shows tremendous promise in applications, it faces significant challenges in motion consistency, computational efficiency, and ethical considerations. This survey provides a comprehensive review of diffusion-based video generation, examining its evolution, technical foundations, and practical applications. We present a systematic taxonomy of current methodologies, analyze architectural innovations and optimization strategies, and investigate applications across low-level vision tasks such as denoising and super-resolution. Additionally, we explore the synergies between diffusion-based video generation and related domains, including video representation learning, question answering, and retrieval. Compared to the existing surveys (Lei et al., 2024a; b; Melnik et al., 2024; Cao et al., 2023; Xing et al., 2024c) which focus on specific aspects of video generation, such as human video synthesis (Lei et al., 2024a) or long-form content generation (Lei et al., 2024b), our work provides a broader, more updated, and more fine-grained perspective on diffusion-based approaches with a special section for evaluation metrics, industry solutions, and training engineering techniques in video generation. This survey serves as a foundational resource for researchers and practitioners working at the intersection of diffusion models and video generation, providing insights into both the theoretical frameworks and practical implementations that drive this rapidly evolving field. A structured list of related works involved in this survey is also available on GitHub: https://github.com/Eyeline-Labs/Survey-Video-Diffusion.
@@ 1 Introduction
The rest of the survey is organized as follows: Section˜2 lays the groundwork by exploring foundational concepts, including video generation paradigms such as GAN-based models, auto-regressive models, and diffusion models. Section˜3 focuses on implementation, addressing practical aspects such as datasets, training engineering techniques, and evaluation metrics, along with benchmarking findings to highlight model performance. Section˜4 covers diverse applications, including conditional generation tasks, enhancement methods such as video denoising, inpainting, interpolation, extrapolation, and super-resolution, as well as personalization, consistency, long-video generation, and emerging 3D-aware diffusion models. Finally, Section˜5 discusses the benefits of diffusion-based video generation to other domains, such as video representation learning, retrieval, QA, and 3D and 4D generation, emphasizing its broader impact in advancing related fields. This structure provides a comprehensive and cohesive understanding of diffusion-based video generation, from foundational principles to advanced applications.
2 Foundations
2.1 Video generative paradigms
@@ 2.1.1 GAN video models
Generative Adversarial Networks (GANs) employ a two-player minimax game between a generator and discriminator, where the generator learns to produce realistic samples while the discriminator distinguishes between generated and real data. The objective function (Goodfellow et al., 2014) is formulated as:
$\min_{G}\max_{D}V(D,G)=\mathbb{E}_{x\sim p_{data}(x)}[\log D(x)]+\mathbb{E}_{z\sim p_{z}(z)}[\log(1-D(G(z)))]$ (1)
where $G$ and $D$ are the generator and discriminator networks, $x$ is a real sample, and $z$ is a random noise vector. $D(x)$ indicates the probability that $x$ is real, while $D(G(z))$ reflects how real the generated sample appears. For video generation, GANs have been extended to model temporal consistency through various architectural innovations and training strategies.
Early video GANs focused on temporal modeling through specialized architectures. TGAN (Saito et al., 2017) employs dual generators: a temporal generator creates motion features while an image generator produces video frames. TGAN-v2 (Saito et al., 2020) improves efficiency through multi-resolution generation using cascaded generator modules. MoCoGAN (Tulyakov et al., 2018) disentangles latent space into motion and content components, predicting motion vectors auto-regressively, while DVD-GAN (Clark et al., 2019) generates all frames in parallel with class conditioning. MoCoGAN-HD (Tian et al., 2020) advances motion-content disentanglement through latent motion trajectory prediction. DIGAN (Yu et al., 2021) leverages implicit neural representations for efficient long-sequence generation with frame-pair focused motion discrimination.
StyleGAN2 (Karras et al., 2020) introduced expressive latent spaces for photorealistic synthesis, establishing foundations for video generation through sparse motion cues. StyleGAN-V (Skorokhodov et al., 2022) models videos as continuous-time signals, enhancing temporal coherence for high-resolution, long-duration generation. StyleInv (Wang et al., 2023i) proposes non-autoregressive GAN inversion for smoother motion transitions and style transfer. AniFaceGAN (Wu et al., 2022c) adapts StyleGAN2 into 3D-aware GANs for multiview-consistent face animation. Recent advances include SIDGAN (Muaz et al., 2023) for high-resolution dubbed video generation through shift-invariant learning and pyramid-based decoding, and StyleLipSync (Ki & Min, 2023) for temporally consistent lip-synced videos with pose-aware masking and latent smoothing.
Large-scale video generative models have emerged as a significant advancement. Open and Advanced Large-Scale Video Generative Models (Wan et al., 2025) demonstrate the potential of scaling GAN architectures for complex video synthesis tasks. Hunyuan-Large (Sun et al., 2024b) introduces a 52-billion parameter mixture-of-experts model, demonstrating the ability of large-scale architectures to handle complex video generation tasks.These advances go beyond traditional GANs and bring models that are more powerful and efficient for large-scale video generation.
@@ 2.1.3 Video diffusion models
GAN-based and Auto-Regressive models often face challenges in maintaining temporal consistency and computational efficiency when generating videos. While GANs struggle with maintaining temporal consistency and often produce unstable or inconsistent frames, diffusion models excel at generating smooth, coherent sequences, particularly for long videos, because their iterative denoising process ensures frame-to-frame consistency. Also, despite Auto-Regressive model can handle both continuous and discrete data, it normally requires high computational cost. In terms of frame quality, diffusion models generate more detailed and higher-resolution outputs compared to GANs, which are prone to artifacts from adversarial training, and autoregressive models, which face resolution limitations due to the compounding of errors in sequential prediction (Moser et al., 2024). Diffusion models also offer superior controllability and local editing capabilities because they can refine specific parts of the video during the denoising process without affecting other regions while GANs and autoregressive methods generate frames in a more holistic or sequential manner (Huang et al., 2024a).
@@ 2.1.4 Auto-regressive video diffusion models
Some recent works further exploit auto-regressive framework in the context of video generation with diffusion models (Weng et al., 2024; Jin et al., 2024; Xie et al., 2024a). ART-V (Weng et al., 2024) first utilized autoregressive diffusion to generate frames sequentially, employing masked diffusion to enhance coherence and reduce inconsistencies. Subsequently, Pyramidal Flow Matching  (Jin et al., 2024) introduced spatial and temporal pyramids to optimize autoregressive generation, reducing computational costs and improving scalability. Progressive Autoregressive Video Diffusion Models  (Xie et al., 2024a) refined this by using a progressive noise schedule, enabling smoother transitions in long videos. CausVid  (Yin et al., 2024) advanced the field by converting bidirectional diffusion models into causal autoregressive generators through asymmetric distillation, achieving real-time frame-by-frame generation with low latency and reduced error accumulation.
2.2 Learning foundations
Video generation relies on different methods to produce realistic and coherent results. As the section discussed above, common paradigms in Figure 1 include GANs, autoregressive models, and diffusion models, each with trade-offs in quality and efficiency. This section introduces the core learning principles behind these methods, starting with diffusion processes, which have become the leading approach for high-quality and temporally consistent video generation. These foundational concepts serve as the basis for the specific diffusion model frameworks and implementations discussed in subsequent sections.
@@ 2.2.7 Training-free methods
Text2Video-Zero (Khachatryan et al., 2023) introduced cross-frame attention based on pretrained text-to-image diffusion model. FateZero (Qi et al., 2023) enhanced this through intermediate attention maps, and Free-Bloom (Huang et al., 2023a) combined GPT-3 with image diffusion models for improved semantic coherence. PEEKABOO (Jain et al., 2024b) enabled spatio-temporal control through masked diffusion using a latent diffusion model-based video generation model, while ControlVideo (Zhang et al., 2023e) extended ControlNet with an interleaved-frame smoother to generate consistent controllable and structurally smooth video. DreamVideo-2 (Wei et al., 2024b) proposed zero-shot subject-driven video customization with precise motion control, enabling subject-preserving and motion-specific video editing. Following this, VideoElevator (Zhang et al., 2024g) encapsulates T2V to enhance temporal consistency and harnesses T2I to provide more faithful details. Recent innovations include GPT4Motion (Lv et al., 2024)’s combination of GPT-4, Blender, and text-to-image diffusion models for physics-aware generation, AnyV2V (Ku et al., 2024)’s universal editing framework built on an image editing model combined with an existing image-to-video generation model to generate the edited video through temporal feature injection. Video Custom Diffusion (Magic-Me) (Ma et al., 2024e) built a training-free 3D Gaussian Noise Prior for video frame initialization, reconstructing inter-frame correlation to achieve identity-preserved generation.
@@ 2.2.8 Token learning
The learning foundations discussed in this section provide the core principles behind modern video generation. These fundamental concepts provide the mathematical and conceptual paradigm, as shown in Figure 1, which are necessary for understanding the specific diffusion model implementations and architectural designs explored in the following sections. Building upon these foundations, subsection 2.4 examines the practical frameworks and implementations to use these theoretical principles into real-world video generation systems.
2.3 Guidances
@@ 2.3.1 Classifier guidance
Algorithm for classifier-guided DDPM sampling. In the DDPM framework, classifier guidance can be incorporated by modifying the reverse diffusion process. The update at each timestep involves sampling from a Gaussian distribution with a mean adjusted by the classifier gradients. Detailed workflow can be referred to Algorithm algorithm˜1.
Algorithm 1 Classifier-guided DDPM sampling, given a diffusion model $(\mu_{\theta}(x_{t}),\Sigma_{\theta}(x_{t}))$, classifier $p_{\phi}(y|x_{t})$, and gradient scale ss.(Dhariwal & Nichol, 2021)
@@
Algorithm for classifier-guided DDIM sampling. For the DDIM framework, the deterministic nature of the sampling process requires a slightly different approach. The classifier gradient is used to modify the epsilon prediction, which guides the reverse sampling process in DDIM. Deatailed workflow can be referred to Algorithm algorithm˜2.
@@ 2.3.2 Classifier-free guidance
Algorithm for Classifier-Free Guidance describes the sampling procedure for classifier-free guidance, which adjusts the score estimates directly within the diffusion model without requiring any classifier gradients. Algorithms algorithm˜3 and algorithm˜4 describe training and sampling with classifier-free guidance in detail.
@@
3:  $(x,y)\sim p(x,y)$ $\triangleright$ Sample data with conditioning information from the dataset
4:  $y\leftarrow\emptyset$ with probability $p_{\mathrm{uncond}}$ $\triangleright$ Randomly discard conditioning to train unconditionally
5:  $\lambda\sim p(\lambda)$ $\triangleright$ Sample log SNR value
@@
7:  $x_{\lambda}=\alpha_{\lambda}x+\sigma_{\lambda}\epsilon$ $\triangleright$ Corrupt data to the sampled log SNR value
8:  Take gradient step on $\nabla_{\theta}\|\epsilon_{\theta}(x_{\lambda},y)-\epsilon\|^{2}$ $\triangleright$ Optimize the denoising model
@@
[…]
@@
6:  $\tilde{\epsilon}_{t-1}\leftarrow(1+w)\epsilon_{\theta}(x_{t-1})-w\epsilon_{\theta}(x_{t-1})$ $\triangleright$ Form the classifier-free guided score at log SNR $\lambda_{t-1}$
7:  $x_{t}\leftarrow\frac{(x_{t-1}-\alpha_{\lambda_{t-1}}\tilde{\epsilon}_{t-1})}{\alpha_{\lambda_{t-1}}}$ $\triangleright$ Sampling step (could be replaced by another sampler, e.g., DDIM)
@@ 2.4 Diffusion model frameworks
This section builds on the theoretical foundations from subsection 2.2 by introducing practical diffusion model frameworks for video generation. While the previous section focused on core principles, this section focus on illustrating how diffusion-based video generation builds on theoretical foundations through a range of practical strategies. The following sections examine four key implementation approaches: pixel-based and latent-based modeling, optical flow integration, adaptive noise scheduling, and agent-driven frameworks, each addressing specific challenges in video synthesis such as temporal consistency, efficiency, and controllability.
@@ 2.4.1 Pixel diffusion and latent diffusion
The theoretical diffusion processes introduced in subsection 2.2 can be implemented through different architectural approaches, each offering distinct advantages for video generation. Pixel-based and latent-based diffusion models represent two fundamental implementation strategies that directly translate the mathematical paradigm of DDPM, DDIM, and other diffusion variants into practical video generation systems as shown in Figure 1.
@@ 2.4.2 Optical-flow-based diffusion models
Another distinct approach within diffusion models leverages optical flow to maintain temporal coherence between frames. Optical flow is a common approach to represent motion by estimating the displacement field between consecutive frames. Early approaches treated it as an optimization problem using handcrafted features to maximize visual similarity(Horn & Schunck, 1981; Black & Anandan, 1993; Bruhn et al., 2005; Sun et al., 2014). Incorporating deep learning-based methods revolutionized this field. FlowNet(Dosovitskiy et al., 2015) pioneered end-to-end optical flow estimation, demonstrating the potential of deep learning in this domain. Subsequent advancements on improved architectures and synthetic datasets further promoted the progress of optical flow estimation(Ilg et al., 2017; Ranjan & Black, 2017; Sun et al., 2018; 2021; Hui et al., 2018; 2020; Yang & Ramanan, 2019). RAFT(Teed & Deng, 2020) included iterative refinement with correlation volumes, significantly boosting performance, while FlowFormer(Huang et al., 2022b; Shi et al., 2023c) applied attention mechanisms within optical flow usage. VideoFlow(Shi et al., 2023b) extended optical flow to video generation scenarios and utilized temporal information across multiple frames, achieving high accuracy. (Hu et al., 2023)propose a Dynamic Multi-scale Voxel Flow Network (DMVFN) to explicitly model the complex motion cues of diverse scales between adjacent video frames by dynamic optical flow estimation to generate videos at lower computational costs. Also, Motion-I2V(Shi et al., 2024) further refined the motion modelling by a diffusion-based motion field predictor and motion-augmented temporal attention to generate more consistent videos even in the presence of large motion and viewpoint variation. By conditioning the optical flow between a previous and a subsequent frame, this framework achieves realistic motion continuity, which is critical for applications in video generation and motion synthesis. Additionally, recent advancements explore upgrading pretrained image diffusion models to video generation by temporally warping input noise. How I Warped Your Noise (Chang et al., 2024a) proposed a novel noise representation called integral noise ($\int$-noise), preserving temporal correlations through an optical-flow-based noise warping algorithm. This foundational approach mitigated issues like high-frequency flickering and texture-sticking artifacts, laying the groundwork for temporally coherent video diffusion. Building upon this, Infinite-Resolution Integral Noise Warping (Deng et al., 2025) method introduces a training-free approach where temporal correlations in noise are preserved by warping input noise using integral representations, enabling consistent frame generation while reducing computational overhead. These diffusion model frameworks demonstrate the versatility and adaptability of diffusion processes across different domains, from high-resolution static images to complex, temporally coherent videos. As shown in Figure 3, Go-with-the-Flow (Burgert et al., 2025) combined optical flow extraction from training videos with real-time noise warping and fine-tuning video diffusion models on paired warped noise and video data to enable robust and motion-controllable generation. Together, these methods demonstrate the progression from temporal correlation preservation to advanced motion control, showcasing the versatility and adaptability of diffusion models in video generation.